OpenStudy (anonymous):

1. Find the 40th term of the sequence 2, -12, 72, ... 2. What is the sum of the first 30 terms of the series 9, 3, 1, ...?

OpenStudy (ddcamp):

40th term? That's a REALLY big number...

OpenStudy (imstuck):

You need to know the formula for geometric sequence with this one. Do you know it?

OpenStudy (imstuck):

It is\[a _{n}=a _{1}(r)^{n-1}\]where n is the term number you are looking for (here it's 40), a1 is the first number in the series (here it's 2), and r is the common ratio (here it's what you are doing to each number to get to the next one. Here we are multiplying each term by -6. Check it to make sure you understand.). Then fill in what you know!

OpenStudy (imstuck):

Can you do this or do you need assistance?

OpenStudy (imstuck):

DDCamp is right...it's a HUGE number.

OpenStudy (imstuck):

Actually, no I take that back, it's incredibly SMALL!!!! Cuz it's negative!

OpenStudy (anonymous):

is it -4.456^30???

OpenStudy (imstuck):

Yes, it is! did you know that already or did the formula help? Jw.

OpenStudy (imstuck):

What about the next one, the sum of the geometric series?

OpenStudy (anonymous):

The formula really helped! Could you possibly help me with the second one? I'm really stuck on that one.

OpenStudy (anonymous):

So i got the sequence is dividing by 3 each term, and now I'm just really confused.

OpenStudy (imstuck):

Looking right now, ok?\[S _{n}=\frac{ a _{1}(1-r ^{n}) }{ 1-r }\]The same thing applies here but with a different formula. n is the number of terms (here it's 30), a1 is the first term in the series (here it's 9), and r is the ratio or what do you do to the second number to get it to the first (just like what do you do to the third to get it to the second). Here this one is a bit tricky, because in order to get from a 3 to a 9, you have to divide by something. That's the tricky part. You have to divide 3 by 1/3. When you do that, you get 9. So r here is 1/3. Fill in your equation now and see what comes out! And hope it's right the way I told you to dot it! lol

OpenStudy (anonymous):

So would it be 536870912/7625597484987? or 0.000007.....???

OpenStudy (imstuck):

It would be very large becasue you are dividing by 1/3 which is actually multiplying by 3 when you flip the fraction and change the division sign to a multiplication sign. Its huge.

OpenStudy (anonymous):

Ok, so Im stuck again! haha I'm not sure what I did wrong! lol

OpenStudy (imstuck):

Now I'm stuck, too. Lol. Seriously! Maybe I'm misremembering my formula! Let me look into it. Just a sec.

OpenStudy (jdoe0001):

so.... what's the "common ratio" or the multiplier from term to term?

OpenStudy (anonymous):

divide by 3?

OpenStudy (jdoe0001):

hmm is supposed to be a multiplier ...... no quite is not 1/3 if that's what you meant

OpenStudy (imstuck):

OMG I'm so dumb! I'm trying to find out why my number seems so strange but I am supposed to be looking for the sum of the terms and instead I was trying to find out the 30th term. Ok, duh, here's what you do. \[S _{n}=\frac{ a _{1}(1-r ^{2}) }{ 1-r }\]Filling in our info we get\[S _{n}=\frac{ 9(1-\frac{ 1 }{ 3 }^{2}) }{ 1-\frac{ 1 }{ 3 } }\]When you take 1/3 to the 30th power you get .000000000000004586. Seriously. So when you subtract that from 1 you will get 1. So then you multiply that 1 by the 9 to get\[S _{n}=\frac{ 9 }{ 1-\frac{ 1 }{ 3 } }\]Do the math and get 27/2, which is 13.5

OpenStudy (anonymous):

Ahh thank you so much!!!

OpenStudy (imstuck):

You're very welcome!